in a recent​ year, a poll asked 2962 random adult citizens of a large country how they rated economic conditions. In the​ poll, 26 ​% rated the economy as​ Excellent/Good. A recent media outlet claimed that the percentage of citizens who felt the economy was in​ Excellent/Good shape was 27 ​%. Does the poll support this​ claim?

Check the conditions. Which of the following conditions are​ satisfied? Select all that apply.

A.Less than​ 10% of the population was sampled.

B.The sample is random.

C.The data are independent.

D.There are more than 10​ "successes" and 10​ "failures."
Answer #1

Option A is correct as the population must be 10 times larger than 2962

option B is correct as it is given that 2962 random adult citizens are used in the sample

option C is correct as each adult citizens is random and must be independent.

option D is correct because np = 2962*0.26 = 770.12 > 10 and n(1-p) = 2962*(1-0.26) = 2192 > 10

Hence, all options are correct